# Tracing Galaxy Evolution in the Nearby Universe: The Role of Dark Matter

### v1

@@ IV Comparison of dynamical and stellar mass
Figure 5: Difference between dynamical and stellar mass ($\Delta{\bf logM_{(Dyn-ED)}}$ = $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ - $\log({\bf M_{ED}/{\bf M_{Solar}}})$) as function of dynamical mass for the LTGs samples. Each row corresponds to a specific estimation of dynamical mass ($\mathbf{M_{KS}}$, $\mathbf{M_{KA}}$, $\mathbf{M_{KB}}$, $\mathbf{M_{nS}}$, $\mathbf{M_{IS}}$, $\mathbf{M_{IA}}$, $\mathbf{M_{IB}}$). Each colour and symbol represents a quasi-constant redshift. The difference in redshift between consecutive symbols is approximately 0.01. The mean uncertainty of the difference between $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ and $\log({\bf M_{ED}/{\bf M_{Solar}}})$ is approximately 0.280 dex. The left column of the mosaic contains the full range of redshift, while the right column only contains three specific redshifts: low (black dots $z\sim 0.025$), intermediate (purple squares $z\sim 0.095$), and high (dark green triangles $z\sim 0.165$), the latter with the aim of appreciating more clearly the differences in dynamical and stellar masses due to redshift and also to display the uncertainties for each point.
@@
Figure 6: Mosaic of the difference between dynamical and stellar mass ($\Delta{\bf logM_{(Dyn-ED)}}$ = $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ - $\log({\bf M_{ED}/{\bf M_{Solar}}})$) as a function of dynamical mass, analogous to Fig. 5. In the first column, the quasi-constant redshift trend lines are spaced every 0.02 instead of 0.01, which reduces crowding and makes the overall behavior easier to visualize. In the second column, three representative redshifts are shown: low (light green diamonds $z\sim 0.045$), intermediate (magenta triangles $z\sim 0.085$), and high (brown diamonds $z\sim 0.145$). These examples complement those shown in Fig. 5, reinforcing the robustness of the behaviors discussed and facilitating the identification of completeness ranges shared by the different trends.
@@
Overall, this mosaic in Fig. 5 reveals a more complex behavior than that shown in Fig. 4. For bulge-dominated LTGs (see the case of $\mathbf{M_{KS}}$, $\mathbf{M_{nS}}$, and $\mathbf{M_{IS}}$) we can see that the vast majority of galaxies are in the relatively high dynamical mass regime (log($\mathbf{M_{Dyn}/M_{Solar}})>11$), we can also see that at low redshift, the mass difference declines marginally with dynamical mass, while at high redshift it increases with dynamical mass, with a gradual transition between these regimes. For disk-dominated LTGs (see the case of $\mathbf{M_{KA}}$, $\mathbf{M_{KB}}$, $\mathbf{M_{IA}}$, and $\mathbf{M_{IB}}$) we can see that galaxies cover an wide dynamical mass interval, we also see that in the relatively low dynamical mass regime the mass difference as a function of dynamical mass exhibits a saddle-like shape at low redshift (see black dots $z\sim 0.025$). This feature evolves with redshift, gradually flattening (see purple squares $z\sim 0.095$) and eventually transforming into an almost straight line with a steadily increasing slope (dark green triangles $z\sim 0.165$). In all cases, within the high-mass regime (log($\mathbf{M_{Dyn}/M_{Solar}})>11$), the behavior closely resembles that reported for early-type galaxies (ETGs) by Nigoche-Netro et al. (2016) (see their Figure 5); specifically, the mass difference increases as a function of dynamical mass and decreases as a function of redshift. Both disk- and bulge-dominated LTGs display the saddle-like feature, though it is more pronounced in the disk-dominated systems. In Appendix C, we present mosaics for the remaining stellar mass estimates ($\mathbf{M_{EN}}$, $\mathbf{M_{WD}}$, $\mathbf{M_{WN}}$, $\mathbf{M_{PP}}$, $\mathbf{M_{P03}}$, $\mathbf{M_{P11}}$, $\mathbf{M_{SF}}$), which confirm the same general trends. As before, the variables and possible biases outlined in Appendix A do not significantly alter the observed patterns. The only exceptions are when dynamical masses are derived from stars ($\mathbf{M_{KS}}$, $\mathbf{M_{nS}}$, $\mathbf{M_{IS}}$) or gas ($\mathbf{M_{KA}}$, $\mathbf{M_{KB}}$, $\mathbf{M_{IA}}$, $\mathbf{M_{IB}}$), reinforcing the conclusion that differences are mainly linked to the structural components of galaxies. In summary, while the detailed mass dependence varies between bulge- and disk-dominated systems, the overall evolutionary trend is consistent: at low redshift, the difference follows a saddle-like shape, which with increasing redshift evolves into a straight line with a progressively steeper slope.
@@
In Figure 7, we present the distribution of dynamical mass as a function of redshift with each data point representing a galaxy. It can be observed that the different selection criteria in mass definitions imposed on the original sample of approximately 126,000 LTGs, as described in III.1 and III.3, reduce the redshift range to an approximate interval of 0-0.25. Samples with information on stellar velocity dispersion ($\mathbf{M_{KS}}$, $\mathbf{M_{nS}}$, and $\mathbf{M_{IS}}$) contain approximately 5,000 LTGs, while samples with gas velocity dispersion information ($\mathbf{M_{KA}}$, $\mathbf{M_{KB}}$, $\mathbf{M_{IA}}$ and $\mathbf{M_{IB}}$) contain approximately 15,000 LTGs. The figure 7 unveils the Malmquist bias (see Section A.5(c)) affecting our galaxy samples. This bias becomes evident as we observe that, at greater distances, only the most luminous and massive objects are sampled. At relatively low redshifts ($z\sim 0.025$), the samples demonstrate approximate completeness for $\log(\mathbf{M_{Dyn}/M_{Solar}})>10.5$. Similarly, at intermediate redshifts ($z\sim 0.095$), the samples are approximately complete for $\log(\mathbf{M_{Dyn}/M_{Solar}})>11$, while at relatively high redshifts ($z\sim 0.165$), the samples are approximately complete for $\log(\mathbf{M_{Dyn}/M_{Solar}})>11.5$. Figure 8 shows the distribution of stellar mass as a function of redshift, revealing a behavior similar to that observed in Figure 7. In fact, the completeness limit at each redshift for stellar masses is shifted toward lower masses by approximately 1.0 dex. This information is crucial in affirming the consistency of the trends shown in Figure 5. Specifically, the observed difference in dynamical and stellar mass as a function of dynamical mass and redshift do not appear to stem from biases in sample completeness, because the comparisons are made exclusively within complete mass ranges and within the ranges shared by the different trend lines. A comprehensive analysis of this point is provided in Section V.6.
@@
Figure 7: Dynamical mass as a function of redshift. Each point represents a galaxy. Each graph corresponds to a specific estimation of dynamical mass ($\mathbf{M_{KS}}$, $\mathbf{M_{KA}}$, $\mathbf{M_{KB}}$, $\mathbf{M_{nS}}$, $\mathbf{M_{IS}}$, $\mathbf{M_{IA}}$, $\mathbf{M_{IB}}$).
@@
Figure 8: Stellar mass as a function of redshift. Each point represents a galaxy. Each graph corresponds to a specific estimation of stellar mass ($\mathbf{M_{ED}}$, $\mathbf{M_{EN}}$, $\mathbf{M_{WD}}$, $\mathbf{M_{WN}}$, $\mathbf{M_{PP}}$, $\mathbf{M_{P03}}$, $\mathbf{M_{P11}}$, $\mathbf{M_{SF}}$).
@@
In Figure 6, we present a mosaic with the same structure as Fig. 5, with identical axes and panel organization. However, in the first column of this figure, the quasi-constant-redshift trend lines are spaced every 0.02 in redshift, instead of every 0.01 as in Fig. 5. This modification reduces the crowding of trend lines, leaving more separation between them and thus clarifying the overall behavior. In the second column of Fig. 6, instead of the three redshifts $z\sim 0.025$, $z\sim 0.095$, and $z\sim 0.165$, we display three alternative values that also represent relatively low, intermediate, and high redshifts: light green diamonds ($z\sim 0.045$), magenta triangles ($z\sim 0.085$), and brown diamonds ($z\sim 0.145$). These additional cases allow us to more clearly identify the completeness ranges shared by the different trends and further support the analysis made for Fig. 5. Taken together, Fig. 6 confirms the behaviors discussed for Fig. 5, but with enhanced clarity and with complementary examples of the evolution of the mass difference at representative redshifts.
@@ V Discussion
In the preceding section, we found a dependence of the difference between the dynamical and stellar mass of LTGs on both the dynamical mass and the redshift. Notably, LTGs dominated by the disk behave differently from those dominated by the bulge and the transition from one regime to another seems complex. Next, we will analyze possible biases due to mass estimations. Once the potential biases and uncertainties associated with mass estimations are characterized, we will discuss the statistical significance of the observed behaviors. Additionally, our discussion extends to the role that the SPS models and IMF might play in the observed behavior, considering that the SPS models might underestimate the baryonic mass and the IMF utilized in estimating stellar masses might not be universally applicable.
@@
The uncertainties in the dynamical mass were computed by us taking into account the SDSS uncertainties in each of the parameters involved. Nevertheless, the primary source of uncertainty stems from the calibration process. As described in Section III.2, estimating the dynamical masses of LTGs is particularly challenging due to the significant influence of the galaxies’ inclination angles. This section explains the uncertainties associated with dynamical mass estimations obtained through galaxy mass calibration. According to Nigoche-Netro et al. (2022), a stronger relationship is observed in the calibration of dynamical masses for galaxies with higher inclination angles ($i>66^{o}$). This relationship is primarily attributed to reduced projection effects at higher inclinations, leading to more accurate mass estimates. Nigoche-Netro et al. (2022) report that the dispersion in calibration fits for these galaxies is approximately 0.23-0.25 on a logarithmic scale (see their Table 6), indicating the level of uncertainty in dynamical mass estimates. In the primary analysis of this study, we focus mainly on LTGs with inclination angles greater than $66^{o}$, resulting in dynamical mass estimates that could vary within this dispersion range from the calculated value. The dispersion quantifies this uncertainty, which can be contextualized by evaluating both the absolute and relative uncertainties of the estimated dynamical mass. On a linear scale, this uncertainty is approximately 50%.
@@
It is worth noting that the individual uncertainties in dynamical and stellar mass are displayed in the various graphs of this study only when the figures have relatively few points and lines and the inclusion of error bars did not impede the visualization of the depicted behaviors.
@@
    Stellar masses with significant differences. In the case of stellar masses $\mathbf{M_{ED}}$, $\mathbf{M_{EN}}$, $\mathbf{M_{WD}}$, $\mathbf{M_{WN}}$, and $\mathbf{M_{PP}}$ considering all estimations of dynamical mass, the differences between dynamical and stellar mass as a function of dynamical mass and redshift show clear underlying trends (see Figures 5, C1, C2, C3, C4).
@@
    Stellar masses with mixed significance. Regarding stellar masses $\mathbf{M_{P03}}$, $\mathbf{M_{P11}}$, and $\mathbf{M_{SF}}$ it is found that when considering dynamical masses obtained from stars, the difference between dynamical and stellar mass exhibit considerable dispersion and less distinct underlying trends. However, for the dynamical masses obtained from gas, the trends are more evident (see Figures C5, C6, C7).
@@
In Figure 9, graphs are presented showing the difference between dynamical and stellar mass as a function of dynamical mass for the seven dynamical masses and the eight stellar masses using three different redshifts (low, medium, and high) with the aim of observing possible differences due to the methods used for the estimation of these masses. Here it is important to emphasize that the samples at low ($z\sim 0.025$), medium ($z\sim 0.095$), and high redshifts ($z\sim 0.165$) are approximately complete for $\log({\bf M_{Dyn}/M_{Solar}})$ greater than 10.5, 11, and 11.5, respectively. In the said Figure 9, it can be observed that the general behavior is similar in all cases, as discussed in the previous paragraph. It is noteworthy that when dynamic masses obtained from stars are used (see $\mathbf{M_{KS}}$, $\mathbf{M_{nS}}$, and $\mathbf{M_{IS}}$), there is a relatively high dispersion among the estimations of the different stellar masses, and this dispersion is similar at different redshifts. In the case of dynamical masses obtained using gas (see $\mathbf{M_{KA}}$, $\mathbf{M_{KB}}$, $\mathbf{M_{IA}}$, and $\mathbf{M_{IB}}$), it is observed that at low redshift, the dispersion is relatively lower than at medium and high redshifts. However, it is also important to mention that the dispersion in the difference between dynamical and stellar mass due to different estimations of stellar mass is of the order of or smaller than the associated uncertainties (except for $\mathbf{M_{SF}}$), so that these differences due to the methods used for the estimation of the stellar mass are not significant. It can be seen that the stellar mass $\mathbf{M_{SF}}$ significantly deviates from the rest of the stellar masses. Figure 9 does not include error bars, since the trend lines of the different stellar mass estimation methods (eight in total) are very close to each other, and adding error bars would increase the crowding of the plots. To address this, we include Figure 10, where only three particular trend lines are shown (the cases of $\mathbf{M_{ED}}$, $\mathbf{M_{EN}}$, and $\mathbf{M_{PP}}$), allowing error bars to be displayed more clearly without overlapping with a dense set of curves.
@@
Figure 9: Difference between dynamical and stellar mass ($\Delta{\bf logM_{(Dyn-Stellar)}}$ = $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ - $\log({\bf M_{Stellar}/{\bf M_{Solar}}})$) as function of dynamical mass for the LTGs samples and three particular redshifts (low, medium, and high). Each row corresponds to a specific estimation of dynamical mass ($\mathbf{M_{KS}}$, $\mathbf{M_{KA}}$, $\mathbf{M_{KB}}$, $\mathbf{M_{nS}}$, $\mathbf{M_{IS}}$, $\mathbf{M_{IA}}$, $\mathbf{M_{IB}}$). Each colour and symbol represents a specific stellar mass. The mean uncertainty of the difference between $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ and $\log({\bf M_{Stellar}/{\bf M_{Solar}}})$ is approximately 0.280, 0.279, 0.285, 0.281, 0.284, 0.331, 0.312, and 0.294 dex for $\mathbf{M_{ED}}$, $\mathbf{M_{EN}}$, $\mathbf{M_{WD}}$, $\mathbf{M_{WN}}$, $\mathbf{M_{PP}}$, $\mathbf{M_{P03}}$, $\mathbf{M_{P11}}$, and $\mathbf{M_{SF}}$ respectively. The samples at low ($z\sim 0.025$), medium ($z\sim 0.095$), and high redshifts ($z\sim 0.165$) are approximately complete for $\log({\bf M_{Dyn}/M_{Solar}})$ greater than 10.5, 11, and 11.5, respectively. $\bf<N>$} is the average number of galaxies in the different samples considered.
@@
Figure 10: Difference between dynamical and stellar mass ($\Delta{\bf logM_{(Dyn-Stellar)}}$ = $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ - $\log({\bf M_{Stellar}/{\bf M_{Solar}}})$) as a function of dynamical mass for the LTG samples at three representative redshifts (low, medium, and high), analogous to Fig. 9. Each row corresponds to a specific estimation of dynamical mass ($\mathbf{M_{KS}}$, $\mathbf{M_{KA}}$, $\mathbf{M_{KB}}$, $\mathbf{M_{nS}}$, $\mathbf{M_{IS}}$, $\mathbf{M_{IA}}$, $\mathbf{M_{IB}}$). Each color and symbol corresponds to a specific stellar mass estimate. The mean uncertainties in the difference between $\log({\bf M_{Dyn}/M_{Solar}})$ and $\log({\bf M_{Stellar}/M_{Solar}})$ are 0.280, 0.279, and 0.284 dex for $\mathbf{M_{ED}}$, $\mathbf{M_{ND}}$, and $\mathbf{M_{PP}}$, respectively. The samples are approximately complete for $\log({\bf M_{Dyn}/M_{Solar}})>10.5$, 11.0, and 11.5 at low ($z\sim 0.025$), medium ($z\sim 0.095$), and high ($z\sim 0.165$) redshift, respectively. $\bf<N>$} is the average number of galaxies in the different samples considered.

### v2

@@ IV Comparison of dynamical and stellar mass
Figure 5: Difference between dynamical and stellar mass ($\Delta{\bf logM_{(Dyn-ED)}}$ = $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ - $\log({\bf M_{ED}/{\bf M_{Solar}}})$) as function of the dynamical mass for the LTGs samples. Each row corresponds to a specific estimation of dynamical mass ($\mathbf{M_{KS}}$, $\mathbf{M_{KA}}$, $\mathbf{M_{KB}}$, $\mathbf{M_{nS}}$, $\mathbf{M_{IS}}$, $\mathbf{M_{IA}}$, $\mathbf{M_{IB}}$). Each color and symbol represents a quasi-constant redshift. The difference in redshift between consecutive symbols is approximately 0.01. The mean uncertainty of the difference between $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ and $\log({\bf M_{ED}/{\bf M_{Solar}}})$ is approximately 0.280 dex. The left column of the mosaic contains the full range of redshift, while the right column only contains three specific redshifts: low (black dots, $z\sim 0.025$), intermediate (purple squares, $z\sim 0.095$), and high (dark green triangles, $z\sim 0.165$), the latter with the aim of appreciating more clearly the differences in dynamical and stellar masses due to redshift and also to display the uncertainties for each point.
@@
Figure 6: Mosaic of the difference between dynamical and stellar mass ($\Delta{\bf logM_{(Dyn-ED)}}$ = $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ - $\log({\bf M_{ED}/{\bf M_{Solar}}})$) as a function of the dynamical mass, analogous to Fig. 5. In the first column, the quasi-constant redshift trend lines are spaced every 0.02 instead of 0.01, which reduces crowding and makes the overall behavior easier to visualize. In the second column, three representative redshifts are shown: low (light green diamonds, $z\sim 0.045$), intermediate (magenta triangles, $z\sim 0.085$), and high (brown diamonds, $z\sim 0.145$). These examples complement those shown in Fig. 5, reinforcing the robustness of the behaviors discussed and facilitating the identification of completeness ranges shared by the different trends.
@@
Overall, this mosaic in Fig. 5 reveals a more complex behavior than that shown in Fig. 4. For bulge-dominated LTGs (see the case of $\mathbf{M_{KS}}$, $\mathbf{M_{nS}}$, and $\mathbf{M_{IS}}$) we can see that the vast majority of galaxies are in the relatively high dynamical mass regime (log($\mathbf{M_{Dyn}/M_{Solar}})>11$), we can also see that at low redshift, the mass difference declines marginally with dynamical mass, while at high redshift it increases with dynamical mass, with a gradual transition between these regimes. For disk-dominated LTGs (see the case of $\mathbf{M_{KA}}$, $\mathbf{M_{KB}}$, $\mathbf{M_{IA}}$, and $\mathbf{M_{IB}}$) we can see that galaxies cover a wide dynamical mass interval. We also see that in the relatively low dynamical mass regime, the mass difference as a function of the dynamical mass exhibits a saddlelike shape at low redshift (see black dots, $z\sim 0.025$). This feature evolves with redshift, gradually flattening (see purple squares, $z\sim 0.095$) and eventually transforming into an almost straight line with a steadily increasing slope (dark green triangles, $z\sim 0.165$). In all cases, within the high-mass regime (log($\mathbf{M_{Dyn}/M_{Solar}})>11$), the behavior closely resembles that reported for ETGs by Nigoche-Netro et al. (2016, see their Figure 5); in particular, the mass difference increases as a function of the dynamical mass and decreases as a function of the redshift. Both disk- and bulge-dominated LTGs display the saddlelike feature, though it is more pronounced in the disk-dominated systems. In Appendix C, we present mosaics for the remaining stellar mass estimates ($\mathbf{M_{EN}}$, $\mathbf{M_{WD}}$, $\mathbf{M_{WN}}$, $\mathbf{M_{PP}}$, $\mathbf{M_{P03}}$, $\mathbf{M_{P11}}$, $\mathbf{M_{SF}}$), which confirm the same general trends. As before, the variables and possible biases outlined in Appendix A do not significantly alter the observed patterns. The only exceptions are when dynamical masses are derived from stars ($\mathbf{M_{KS}}$, $\mathbf{M_{nS}}$, $\mathbf{M_{IS}}$) or gas ($\mathbf{M_{KA}}$, $\mathbf{M_{KB}}$, $\mathbf{M_{IA}}$, $\mathbf{M_{IB}}$), reinforcing the conclusion that differences are mainly linked to the structural components of galaxies. In summary, while the detailed mass dependence varies between bulge- and disk-dominated systems, the overall evolutionary trend is consistent; at low redshift, the difference follows a saddlelike shape, which with increasing redshift evolves into a straight line with a progressively steeper slope.
@@
In Figure 7, we present the distribution of dynamical mass as a function of the redshift, with each data point representing a galaxy. It can be observed that the different selection criteria in mass definitions imposed on the original sample of approximately 126,000 LTGs, as described in III.1 and III.3, reduce the redshift range to an approximate interval of 0–0.25. Samples with information on stellar velocity dispersion ($\mathbf{M_{KS}}$, $\mathbf{M_{nS}}$, and $\mathbf{M_{IS}}$) contain approximately 5000 LTGs, while samples with gas velocity dispersion information ($\mathbf{M_{KA}}$, $\mathbf{M_{KB}}$, $\mathbf{M_{IA}}$ and $\mathbf{M_{IB}}$) contain approximately 15,000 LTGs. The figure 7 unveils the Malmquist bias (see Appendix A.5(c)) affecting our galaxy samples. This bias becomes evident as we observe that, at greater distances, only the most luminous and massive objects are sampled. At relatively low redshifts ($z\sim 0.025$), the samples demonstrate approximate completeness for $\log(\mathbf{M_{Dyn}/M_{Solar}})>10.5$. Similarly, at intermediate redshifts ($z\sim 0.095$), the samples are approximately complete for $\log(\mathbf{M_{Dyn}/M_{Solar}})>11$, while at relatively high redshifts ($z\sim 0.165$), the samples are approximately complete for $\log(\mathbf{M_{Dyn}/M_{Solar}})>11.5$. Figure 8 shows the distribution of stellar mass as a function of the redshift, revealing a behavior similar to that observed in Figure 7. In fact, the completeness limit at each redshift for stellar masses is shifted toward lower masses by approximately 1.0 dex. This information is crucial in affirming the consistency of the trends shown in Figure 5. In particular, the observed difference in the dynamical and stellar mass as a function of dynamical mass and redshift does not appear to stem from biases in sample completeness, because the comparisons are made exclusively within complete mass ranges and within the ranges shared by the different trend lines. A comprehensive analysis of this point is provided in Section V.6.
@@
Figure 7: Dynamical mass as a function of the redshift. Each point represents a galaxy. Each graph corresponds to a specific estimation of dynamical mass ($\mathbf{M_{KS}}$, $\mathbf{M_{KA}}$, $\mathbf{M_{KB}}$, $\mathbf{M_{nS}}$, $\mathbf{M_{IS}}$, $\mathbf{M_{IA}}$, $\mathbf{M_{IB}}$).
@@
Figure 8: Stellar mass as a function of the redshift. Each point represents a galaxy. Each graph corresponds to a specific estimation of stellar mass ($\mathbf{M_{ED}}$, $\mathbf{M_{EN}}$, $\mathbf{M_{WD}}$, $\mathbf{M_{WN}}$, $\mathbf{M_{PP}}$, $\mathbf{M_{P03}}$, $\mathbf{M_{P11}}$, $\mathbf{M_{SF}}$).
@@
In Figure 6, we present a mosaic with the same structure as Fig. 5, with identical axes and panel organization. However, in the first column of this figure, the quasi-constant-redshift trend lines are spaced every 0.02 in redshift, instead of every 0.01 as in Fig. 5. This modification reduces the crowding of trend lines, leaving more separation between them and thus clarifying the overall behavior. In the second column of Fig. 6, instead of the three redshifts, $z\sim 0.025$, $z\sim 0.095$, and $z\sim 0.165$, we display three alternative values that also represent relatively low, intermediate, and high redshifts: light green diamonds ($z\sim 0.045$), magenta triangles ($z\sim 0.085$), and brown diamonds ($z\sim 0.145$). These additional cases allow us to more clearly identify the completeness ranges shared by the different trends and further support the analysis made for Fig. 5. Taken together, Fig. 6 confirms the behaviors discussed for Fig. 5, but with enhanced clarity and with complementary examples of the evolution of the mass difference at representative redshifts.
@@ V Discussion
In the preceding section, we found a dependence of the difference between the dynamical and stellar mass of LTGs on both the dynamical mass and the redshift. Notably, LTGs dominated by the disk behave differently from those dominated by the bulge, and the transition from one regime to another seems complex. Next, we will analyze possible biases due to mass estimations. Once the potential biases and uncertainties associated with mass estimations are characterized, we will discuss the statistical significance of the observed behaviors. Additionally, our discussion extends to the role that the SPS models and IMF might play in the observed behavior, considering that the SPS models might underestimate the baryonic mass, and the IMF utilized in estimating stellar masses might not be universally applicable.
@@
The uncertainties in the dynamical mass were computed by us, taking into account the SDSS uncertainties in each of the parameters involved. Nevertheless, the primary source of uncertainty stems from the calibration process. As described in Section III.2, estimating the dynamical masses of LTGs is particularly challenging due to the significant influence of the galaxies’ inclination angles. This section explains the uncertainties associated with dynamical mass estimations obtained through galaxy mass calibration. According to Nigoche-Netro et al. (2022), a stronger relationship is observed in the calibration of dynamical masses for galaxies with higher inclination angles ($i>66^{o}$). This relationship is primarily attributed to reduced projection effects at higher inclinations, leading to more accurate mass estimates. Nigoche-Netro et al. (2022) report that the dispersion in calibration fits for these galaxies is approximately 0.23-0.25 on a logarithmic scale (see their Table 6), indicating the level of uncertainty in dynamical mass estimates. In the primary analysis of this study, we focus mainly on LTGs with inclination angles greater than $66^{o}$, resulting in dynamical mass estimates that could vary within this dispersion range from the calculated value. The dispersion quantifies this uncertainty, which can be contextualized by evaluating both the absolute and relative uncertainties of the estimated dynamical mass. On a linear scale, this uncertainty is approximately 50%.
@@
It is worth noting that the individual uncertainties in dynamical and stellar mass are displayed in the various graphs of this study only when the figures have relatively few points and lines, and the inclusion of error bars did not impede the visualization of the depicted behaviors.
@@
    Stellar masses with significant differences. In the case of stellar masses $\mathbf{M_{ED}}$, $\mathbf{M_{EN}}$, $\mathbf{M_{WD}}$, $\mathbf{M_{WN}}$, and $\mathbf{M_{PP}}$ considering all estimations of dynamical mass, the differences between dynamical and stellar mass as a function of the dynamical mass and redshift show clear underlying trends (see Figures 5, C1, C2, C3, C4).
@@
    Stellar masses with mixed significance. Regarding stellar masses $\mathbf{M_{P03}}$, $\mathbf{M_{P11}}$, and $\mathbf{M_{SF}}$ it is found that when considering dynamical masses obtained from stars, the difference between dynamical and stellar mass exhibits considerable dispersion and less distinct underlying trends. However, for the dynamical masses obtained from gas, the trends are more evident (see Figures C5, C6, C7).
@@
In Figure 9, graphs are presented showing the difference between dynamical and stellar mass as a function of the dynamical mass for the seven dynamical masses and the eight stellar masses using three different redshifts (low, medium, and high) with the aim of observing possible differences due to the methods used for the estimation of these masses. Here, it is important to emphasize that the samples at low ($z\sim 0.025$), medium ($z\sim 0.095$), and high redshifts ($z\sim 0.165$) are approximately complete for $\log({\bf M_{Dyn}/M_{Solar}})$ greater than 10.5, 11, and 11.5, respectively. In the said Figure 9, it can be observed that the general behavior is similar in all cases, as discussed in the previous paragraph. It is noteworthy that when dynamic masses obtained from stars are used (see $\mathbf{M_{KS}}$, $\mathbf{M_{nS}}$, and $\mathbf{M_{IS}}$), there is a relatively high dispersion among the estimations of the different stellar masses, and this dispersion is similar at different redshifts. In the case of dynamical masses obtained using gas (see $\mathbf{M_{KA}}$, $\mathbf{M_{KB}}$, $\mathbf{M_{IA}}$, and $\mathbf{M_{IB}}$), it is observed that at low redshift, the dispersion is relatively lower than at medium and high redshifts. However, it is also important to mention that the dispersion in the difference between dynamical and stellar mass due to different estimations of stellar mass is of the order of or smaller than the associated uncertainties (except for $\mathbf{M_{SF}}$), so that these differences due to the methods used for the estimation of the stellar mass are not significant. It can be seen that the stellar mass $\mathbf{M_{SF}}$ significantly deviates from the rest of the stellar masses. Figure 9 does not include error bars, since the trend lines of the different stellar mass estimation methods (eight in total) are very close to each other, and adding error bars would increase the crowding of the plots. To address this, we include Figure 10, where only three particular trend lines are shown (the cases of $\mathbf{M_{ED}}$, $\mathbf{M_{EN}}$, and $\mathbf{M_{PP}}$), allowing error bars to be displayed more clearly without overlapping with a dense set of curves.
@@
Figure 9: Difference between dynamical and stellar mass ($\Delta{\bf logM_{(Dyn-Stellar)}}$ = $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ - $\log({\bf M_{Stellar}/{\bf M_{Solar}}})$) as a function of the dynamical mass for the LTGs samples and three particular redshifts (low, medium, and high). Each row corresponds to a specific estimation of dynamical mass ($\mathbf{M_{KS}}$, $\mathbf{M_{KA}}$, $\mathbf{M_{KB}}$, $\mathbf{M_{nS}}$, $\mathbf{M_{IS}}$, $\mathbf{M_{IA}}$, $\mathbf{M_{IB}}$). Each color and symbol represents a specific stellar mass. The mean uncertainty of the difference between $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ and $\log({\bf M_{Stellar}/{\bf M_{Solar}}})$ is approximately 0.280, 0.279, 0.285, 0.281, 0.284, 0.331, 0.312, and 0.294 dex for $\mathbf{M_{ED}}$, $\mathbf{M_{EN}}$, $\mathbf{M_{WD}}$, $\mathbf{M_{WN}}$, $\mathbf{M_{PP}}$, $\mathbf{M_{P03}}$, $\mathbf{M_{P11}}$, and $\mathbf{M_{SF}}$ respectively. The samples at low ($z\sim 0.025$), medium ($z\sim 0.095$), and high redshifts ($z\sim 0.165$) are approximately complete for $\log({\bf M_{Dyn}/M_{Solar}})$ greater than 10.5, 11, and 11.5, respectively. <N><N> is the average number of galaxies in the different samples considered.
@@
Figure 10: Difference between dynamical and stellar mass ($\Delta{\bf logM_{(Dyn-Stellar)}}$ = $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ - $\log({\bf M_{Stellar}/{\bf M_{Solar}}})$) as a function of the dynamical mass for the LTG samples at three representative redshifts (low, medium, and high), analogous to Fig. 9. Each row corresponds to a specific estimation of dynamical mass ($\mathbf{M_{KS}}$, $\mathbf{M_{KA}}$, $\mathbf{M_{KB}}$, $\mathbf{M_{nS}}$, $\mathbf{M_{IS}}$, $\mathbf{M_{IA}}$, $\mathbf{M_{IB}}$). Each color and symbol corresponds to a specific stellar mass estimate. The mean uncertainties in the difference between $\log({\bf M_{Dyn}/M_{Solar}})$ and $\log({\bf M_{Stellar}/M_{Solar}})$ are 0.280, 0.279, and 0.284 dex for $\mathbf{M_{ED}}$, $\mathbf{M_{ND}}$, and $\mathbf{M_{PP}}$, respectively. The samples are approximately complete for $\log({\bf M_{Dyn}/M_{Solar}})>10.5$, 11.0, and 11.5 at low ($z\sim 0.025$), medium ($z\sim 0.095$), and high ($z\sim 0.165$) redshift, respectively. <N><N> is the average number of galaxies in the different samples considered.